# Electrically switchable non-relativistic Zeeman spin splittings in collinear antiferromagnets

### v1

@@ Abstract
Materials with electric-field switchable NRZSSs — Assisted by the MAGNDATA database [43, 44], and guided by Tables 1 and 2, we explore materials with linear magnetoelectric NRZSSs or bilinear piezomagnetoelectric NRZSSs. We search for nonpolar collinear antiferromagnets composed of light elements  [^5] The strength of SOI basically depends on the atomic mass. Materials composed of heavy elements (e.g., Bi, Sb, and Pt) likely showcase sizable SOI [50, 51, 52]., identify the SPGs for such antiferromagnets, and discover LiMnPO4 and NaMnP as two representatives with electric-field switchable NRZSSs. The antiferromagnetic LiMnPO4 has a Néel temperature of 34 K [46, 47] and an SPG of ${}^{\bar{1}}m^{1}m^{1}m^{\infty m}1$. According to Table 1, the ${}^{\bar{1}}m^{1}m^{1}m^{\infty m}1$ SPG enables linear magnetoelectric NRZSSs via the $E_{x}\sigma_{\chi}$ coupling. The NaMnP material is a room-temperature antiferromagnet with a Néel temperature of 560 K [48, 49]. The SPG of NaMnP is ${}^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m^{\infty m}1$, and this group hosts the piezomagnetoelectric NRZSSs induced by the combination of $\eta_{xy}$ and $E_{z}$ (see the $\eta_{xy}E_{z}\sigma_{\chi}$ coupling in Table 2). To validate our analysis, we compute the band structures for LiMnPO4 and NaMnP antiferromagnets. First of all, we neglect the SOI and carry out collinear magnetism calculations. This treats the magnetic moments of Mn in LiMnPO4 and NaMnP as scalar quantities (see Fig. 1). On this condition, the electronic spin degree of freedom appears as unmixed spin-up $+s$ and spin-down $-s$ states, where $\chi$ in $s_{\chi}$ is omitted at the collinear magnetism level.
@@
Figure 1: Collinear magnetic structures for (a) LiMnPO4 and (b) NaMnP antiferromagnets, where only the Mn sublattices are displayed. The cyan and pink spheres denote Mn ions with magnetic moments along $+\bm{\chi}$ and $-\bm{\chi}$ directions, respectively. In panels (a) and (b), the boxes enclosed by the black dashed lines represent the cells of LiMnPO4 and NaMnP employed in our simulations. The primitive cell of NaMnP is sketched by the box with pink solid lines in panel (b).
@@
As shown in Fig. S2 of the SM, both the conduction band minimum (CBM) of LiMnPO4 and the valence band maximum (VBM) of NaMnP locate at the Brillouin zone center. We shall focus on such two band edges in our following discussion. Under null electric field, the spin-up $\varepsilon_{\mathrm{C}}(+s)$ and spin-down $\varepsilon_{\mathrm{C}}(-s)$ energy levels of LiMnPO4 are degenerate at the CBM. Applying an electric field along the $x$ direction splits the $\varepsilon_{\mathrm{C}}(+s)$ and $\varepsilon_{\mathrm{C}}(-s)$ energy levels. To be accurate, the $\Delta_{1}=\varepsilon_{\mathrm{C}}(+s)-\varepsilon_{\mathrm{C}}(-s)$ splittings are $\mp 23$ meV for $E_{x}=\pm 6$ MV/cm [see Figs. 2(a) and 2(b)]. For unstrained NaMnP under null electric field, the spin-up $\varepsilon_{\mathrm{V}}(+s)$ and spin-down $\varepsilon_{\mathrm{V}}(-s)$ energy levels at the VBM are degenerated as well. In the presence of epitaxial strain $\eta_{xy}=4\%$, the $\Delta_{2}=\varepsilon_{\mathrm{V}}(+s)-\varepsilon_{\mathrm{V}}(-s)$ splittings of ±$\pm 26$ meV can be created by applying $E_{z}$ of ±$\pm 6$ MV/cm. Figure 3 shows the $\Delta_{1}$ splittings in LiMnPO4 as a function of $E_{x}$ and the $\Delta_{2}$ splittings in NaMnP as a function of $E_{z}$ ($\eta_{xy}=4\%$). This confirms that both the magnetoelectric NRZSSs in LiMnPO4 (CBM) and piezomagnetoelectric NRZSSs (VBM) in NaMnP depend linearly on the corresponding electric fields, and are electrically switchable.
@@
To complete this section, we check the relativistic effect on the electric-field induced Zeeman-type spin splittings in LiMnPO4 and NaMnP. We calculate the band structures for LiMnPO4 and NaMnP at noncollinear magnetism level (with SOI). This treats the magnetic moments of Mn as vectors along ±$\pm\bm{\chi}$ directions ($\bm{\chi}=\mathbf{x},\mathbf{y},\mathbf{z}$). Because of the SOI, the spin-up and spin-down eigen states are not well defined in LiMnPO4 and NaMnP. Instead, the electronic spin degree of freedom is characterized by the spin angular momentum vector, whose predominant component is along $+\bm{\chi}$ or $-\bm{\chi}$ direction. In this case, we may use $\Delta^{\prime}_{1}=\varepsilon_{C}(+s_{\chi})-\varepsilon_{C}(-s_{\chi})$ and $\Delta^{\prime}_{2}=\varepsilon_{V}(+s_{\chi})-\varepsilon_{V}(-s_{\chi})$ to describe the Zeeman-type spin splittings associated with the CBM of LiMnPO4 and the VBM of NaMnP, respectively (see the previous paragraph for the definitions of $\Delta_{1}$ and $\Delta_{2}$). In Fig. S3 of the SM, we show the band structures with Zeeman-type spin splittings for LiMnPO4 and NaMnP, calculated at the relativistic level. Basically, the inclusion of relativistic corrections does not affect the magnitude of the Zeeman-type spin splittings in LiMnPO4 and NaMnP.
@@
Figure 2: The electric-field switchable NRZSSs in LiMnPO4 and NaMnP. Panels (a) and (b) are magnetoelectric NRZSSs in LiMnPO4 driven by $E_{x}$, while panels (c) and (d) are piezomagnetoelectric NRZSSs in NaMnP driven by $E_{z}$ combined with $\eta_{xy}=4\%$. The red (blue) solid lines correspond to the spin-up (spin-down) states, and $\varepsilon_{F}$ denotes the fermi level.
@@
Figure 3: The linear magnetoelectric $\Delta_{1}$ splittings in LiMnPO4 and bilinear piezomagnetoelectric $\Delta_{2}$ splittings in NaMnP. Purple circles show $\Delta_{1}$ in LiMnPO4 as a function of $E_{x}$; Blue squares show $\Delta_{2}$ in NaMnP as a function of $E_{z}$ (under a fixed strain of $\eta_{xy}=4\%$). Purple and blue lines display the corresponding linear regression results.
@@
Summary and outlook — By analyzing the symmetries of spin point groups, we develop a theory that describes the NRZSSs in collinear antiferromagnets driven by electric field. In particular, we identify the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms, both of which are associated with NRZSSs depending linearly on the applied electric field. The reversal of electric field flips the electronic spin states, yielding the electric-field switchable NRZSSs. Our central results are summarized in Tables 1 and 2, and this guides the first-principles predictions of linear magnetoelectric NRZSSs in LiMnPO4 and bilinear piezomagnetoelectric NRZSSs in NaMnP.

### v2

@@ Abstract
Following the analysis in Section V of SM, the FeOOH material has an SPG of $^{\bar{1}}m^{1}m^{1}m^{\infty m}1$. According to Table 1, the $^{\bar{1}}m^{1}m^{1}m^{\infty m}1$ SPG enables linear magnetoelectric NRZSSs via the $E_{x}\sigma_{\chi}$ coupling. The SPG of NaMnP is $^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m^{\infty m}1$, and this group hosts the piezomagnetoelectric NRZSSs induced by the combination of $\eta_{xy}$ and $E_{z}$ (see the $\eta_{xy}E_{z}\sigma_{\chi}$ coupling in Table 2). To validate our analysis, we compute the band structures for FeOOH and NaMnP antiferromagnets. First of all, we neglect the SOI and carry out collinear magnetism calculations. This treats the magnetic moments of Fe in FeOOH and Mn in NaMnP as scalar quantities (see Fig. 1). On this condition, the electronic spin degree of freedom appears as unmixed spin-up +s+s and spin-down s-s states, where $\chi$ in $s_{\chi}$ is omitted at the collinear magnetism level.
@@
Figure 1: Collinear magnetic structures for (a) FeOOH and (b) NaMnP antiferromagnets, where only the Fe and Mn sublattices are displayed. The cyan and pink spheres denote ions (Fe or Mn) with magnetic moments along $\bm{\chi}$ and $\bm{\chi}$ directions, respectively. In panels (a) and (b), the boxes enclosed by the black dashed lines represent the cells of FeOOH and NaMnP employed in our simulations. The primitive cell of NaMnP is sketched by the box with pink solid lines in panel (b).
@@
As shown in Fig. S2 of the SM, both the conduction band minimum (CBM) of FeOOH and the valence band maximum (VBM) of NaMnP locate at the Brillouin zone center. We shall focus on such two band edges in our following discussion. Under null electric field, the spin-up $\varepsilon_{\mathrm{C}}(+s)$ and spin-down $\varepsilon_{\mathrm{C}}(-s)$ energy levels of FeOOH are degenerate at the CBM. Applying an electric field along the $x$ direction splits the $\varepsilon_{\mathrm{C}}(+s)$ and $\varepsilon_{\mathrm{C}}(-s)$ energy levels. To be accurate, the $\Delta_{1}=\varepsilon_{\mathrm{C}}(+s)-\varepsilon_{\mathrm{C}}(-s)$ splittings are $\pm 40$ meV for $E_{x}=\pm 6$ MV/cm [see Figs. 2(a) and 2(b)]. For unstrained NaMnP under null electric field, the spin-up $\varepsilon_{\mathrm{V}}(+s)$ and spin-down $\varepsilon_{\mathrm{V}}(-s)$ energy levels at the VBM are degenerated as well. In the presence of epitaxial strain $\eta_{xy}=4\%$, the $\Delta_{2}=\varepsilon_{\mathrm{V}}(+s)-\varepsilon_{\mathrm{V}}(-s)$ splittings of $\pm 26$ meV can be created by applying $E_{z}$ of $\pm 6$ MV/cm. Figure 3 shows the $\Delta_{1}$ splittings in FeOOH as a function of $E_{x}$ and the $\Delta_{2}$ splittings in NaMnP as a function of $E_{z}$ ($\eta_{xy}=4\%$). This confirms that both the magnetoelectric NRZSSs in FeOOH (CBM) and piezomagnetoelectric NRZSSs (VBM) in NaMnP depend linearly on the corresponding electric fields, and are electrically switchable.
@@
To complete this section, we check the relativistic effect on the electric-field induced Zeeman-type spin splittings in FeOOH and NaMnP. We calculate the band structures for FeOOH and NaMnP at noncollinear magnetism level (with SOI). This treats the magnetic moments of Fe and Mn as vectors along $\pm\bm{\chi}$ directions ($\bm{\chi}=\mathbf{x},\mathbf{y},\mathbf{z}$). Because of the SOI, the spin-up and spin-down eigen states are not well defined in FeOOH and NaMnP. Instead, the electronic spin degree of freedom is characterized by the spin angular momentum vector, whose predominant component is along $\bm{\chi}$ or $\bm{\chi}$ direction. In this case, we may use $\Delta^{\prime}_{1}=\varepsilon_{\mathrm{C}}(+s_{\chi})-\varepsilon_{\mathrm{C}}(-s_{\chi})$ and $\Delta^{\prime}_{2}=\varepsilon_{\mathrm{V}}(+s_{\chi})-\varepsilon_{\mathrm{V}}(-s_{\chi})$ to describe the Zeeman-type spin splittings associated with the CBM of FeOOH and the VBM of NaMnP, respectively (see the previous paragraph for the definitions of $\Delta_{1}$ and $\Delta_{2}$). In Fig. S3 of the SM, we show the band structures with Zeeman-type spin splittings for FeOOH and NaMnP, calculated at the relativistic level. Basically, the inclusion of relativistic corrections does not affect the magnitude of the Zeeman-type spin splittings in FeOOH and NaMnP.
@@
Figure 2: The electric-field switchable NRZSSs in FeOOH and NaMnP. Panels (a) and (b) are magnetoelectric NRZSSs in FeOOH driven by $E_{x}$, while panels (c) and (d) are piezomagnetoelectric NRZSSs in NaMnP driven by $E_{z}$ combined with $\eta_{xy}=4\%$. The red (blue) solid lines correspond to the spin-up (spin-down) states, and $\varepsilon_{F}$ denotes the Fermi level.
@@
Figure 3: The linear magnetoelectric $\Delta_{1}$ splittings in FeOOH and bilinear piezomagnetoelectric $\Delta_{2}$ splittings in NaMnP. Purple circles show $\Delta_{1}$ in FeOOH as a function of $E_{x}$; Blue squares show $\Delta_{2}$ in NaMnP as a function of $E_{z}$ (under a fixed strain of $\eta_{xy}=4\%$). Purple and blue lines display the corresponding linear regression results.
@@
Summary and outlook — By analyzing the symmetries of spin point groups, we develop a theory that describes the NRZSSs in collinear antiferromagnets driven by electric field. In particular, we identify the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms, both of which are associated with NRZSSs depending linearly on the applied electric field. The reversal of electric field flips the electronic spin states, yielding the electric-field switchable NRZSSs. Our central results are summarized in Tables 1 and 2, and this guides the first-principles predictions of linear magnetoelectric NRZSSs in FeOOH and bilinear piezomagnetoelectric NRZSSs in NaMnP.